OpenStudy (kitten_is_back):

Which value of x makes the equation true? x – 3.4 = 9.6 x =

OpenStudy (kenialove):

x - 3.4 = 9.6 + 3.4 +3.4 What's 9.6 + 3.4 = ?

OpenStudy (kenialove):

Do you understand ? @Kitten_is_back

OpenStudy (kitten_is_back):

13

OpenStudy (kenialove):

Correct! So x = 13